# Large-order perturbation theory of linear eigenvalue problems

### v1

@@ Abstract
We consider a class of linear eigenvalue problems depending on a small parameter $\epsilon$ in which the series expansion for the eigenvalue in powers of $\epsilon$ is divergent. We develop a new technique to determine the precise nature of this divergence. We illustrate the technique through its application to four examples: the anharmonic oscillator, a simplified model of equitorially-trapped Rossby waves, and two simplified models based on quasinormal modes of Reissner-Normström de Sitterblack holes.
@@ 1 Introduction
Sometimes this series diverges, and it is of interest to determine the nature of this divergence. Here we provide a straightforward approach to determining the precise asymptotic behaviour of $\lambda_{n}$ for large $n$.
@@
The main technique of Bender and Wu is to analytically continue in $\epsilon$ (typically until $\epsilon$ is negative), and then solve the resulting problem by combination of Liouville-Green (WKB) and matched asymptotic approximations. This is a delicate procedure, since the goal is to identify an exponentially small component of $\lambda$ beyond-all-orders of the divergent asymptotic series. Cauchy’s integral formula is then used to determine the coefficients in the power series expansion of $\lambda$ on the positive real $\epsilon$ axis in terms of the values of $\lambda$ on either side of the negative real $\epsilon$ axis. This technique is ingenious, and has proved successful, but the details can be very complicated. The present work aims to present a simpler alternative method, which we hope will be useful.
@@
In the applied mathematics literature an early example of such a problem occurs in the work of Boyd and Natarov [4], who consider a model problem for an equitorially-trapped Rossby wave in a shear flow in the ocean or atmosphere. There the main interest is in the imaginary part of the eigenvalue (corresponding to the growth rate of instability)—the divergent perturbation series is purely real, but there is an exponentially small imaginary part beyond all orders. In [10] this problem is attacked in almost the reverse direction to Bender and Wu—the divergent series is first found, and used to determine the exponentially small imaginary component of the eigenvalue via optimal truncation and Stokes phenomenon, rather than the other way round. Simplifying and extending the procedure from [10] forms the basis of the present work.
@@
We present our procedure through its application to four examples. Each follows the same general framework, which we hope will allow the interested reader to adapt the method to their own particular problem, but the final part of the analysis differs slightly in each case.
@@ 2 Example 1: Simplified black holes
with $g(0)=1$ and $g(x)=o(\mathrm{e}^{-x})$ as $x\rightarrow$-$\infty$, where $0<\epsilon\ll 1$. This is a much-simplified version of the problem in [7] concerning quasinormal modes of Reissner-Normström de Sitter black holes; the eigenvalue $\omega$ is the frequency of the mode.
@@
The boundary conditions at $x=0$ and $x=$-$\infty$ each remove one degree of freedom, so that there is a nonzero solution only if $\omega$ takes particular values. The goal is to find the asymptotic expansion of the leading eigenvalue,
@@
as $\epsilon\rightarrow 0$, and in particular the form of the divergence of $\omega_{n}$ as $n\rightarrow\infty$.
@@ 2.3 Inner region near $X=1$
$c_{n}=-(-1)^{n}\frac{(3/4)_{n-1}(5/4)_{n-1}}{16(2)_{n-1}}.$
Matching with (13) gives
@@ 2.4 Boundary layer in the late terms near $X=0$
is the exponential integal. Now $g_{n}$ should be regular as $\xi\rightarrow 0$. Since $\operatorname{Ei}(2\xi)\sim\log\xi$ as $\xi\rightarrow 0$, we need
@@
to remove the logarithmic singularity at $\xi=0$. Since $\operatorname{Ei}(2\xi)$ exhibits Stokes phenomenon for large $\xi$ there will be a switch in the behaviour of the late terms depending on the argument of $\xi$—this is what is known as the higher-order Stokes phenomenon, a Stokes phenemonon not in the asymptotic expansion of $g$ as a function of $\epsilon$, but in the late-term approximation of $g_{n}$ [6, 9]. There is a higher-order Stokes line when $\xi$ crosses the positive real axis, across which the constant contribution to the large-$\xi$ approximation of $H$ (i.e. in the outer limit of the inner expansion) changes. Note that there is no Stokes phonomenon associated with the particular solution $\log\xi$, so that the coefficient of $\mathrm{e}^{2\xi}/\xi$ is fixed. This will not be the case in our other examples.
@@
Figure 1: Divergence of the coefficients in the asymptotic expansion of $\omega$. (a) coefficients determined numerically from (4)-(5). The linear growth is consistent with factorial divergence. (b) The ratio of the numerical value to the asymptotic prediction (17). Blue is the base series, while orange, green and red correspond to enhanced convergence using Richardson extrapolation on two, three and four terms respectively. The convergence is slower than expected because of the presence of log terms in the higher-order corrections, unaccounted for in the extrapolation.
@@ 3 Example 2: Anharmonic oscillator
Having introduced the procedure with a simple model problem, we now consider the classical problem of the anharmonic oscillator [2, 3]. Much of the analysis follows the same framework, though the details of the boundary layer in the late-term approximation analagous to §2.4 are a little different.
@@ 3.2 Outer region
Substituting (27) into (26) and equating coefficients of powers of $\epsilon$ gives, at leading order, the eikonal equation
$(\phi^{\prime})^{2}+X\phi^{\prime}+\frac{X^{4}}{4}=0.$
The solution which decays at infinity is
$\phi^{\prime}=\frac{X}{2}\left(1-\sqrt{1+X^{2}}\right)$
so that
$\phi=\frac{1}{6}+\frac{X^{2}}{4}-\frac{(1+X^{2})^{3/2}}{6}$
where we have normalised by setting $\phi(0)=0$. The amplitude equation is
$X(1+X^{2})^{1/2}A^{\prime}+\left((1+X^{2})^{1/2}-\frac{1}{2}-\frac{1}{2(1+X^{2})^{1/2}}\right)A-\epsilon A^{\prime\prime}-\epsilon\bar{\lambda}A=0.$
Expanding $A$ in powers of $\epsilon$ as in (27) gives, at leading order,
$A_{0}=\frac{B_{0}}{(1+X^{2})^{1/4}\sqrt{1+\sqrt{1+X^{2}}}},$ (28)
for some constant $B_{0}$. Matching with the inner solution requires $A(0)=1$, so that $B_{0}=\sqrt{2}$. In general, the equation for $A_{n}$ is
@@
As before, there are two sources of divergence in $A_{n}$: the usual factorial/power from repeated differentiation of the singularity in $A_{0}$, and a factorial/constant divergence driven by $\bar{\lambda}_{n}$. For the first, we use the usual ansatz [5]
$A_{n}\sim\frac{G\Gamma(n+\gamma)}{\chi^{n+\gamma}},$
as $n\rightarrow\infty$. At leading order
$X(1+X^{2})^{1/2}-\chi^{\prime}=0$
so that
$\chi=-\frac{(1+X^{2})^{3/2}}{3},$
since $\chi=0$ at $X^{2}=-1$. At next order,
$0$ $=$ $-X(1+X^{2})^{1/2}G^{\prime}+\left(-(1+X^{2})^{1/2}-\frac{1}{2}+\frac{1}{2(1+X^{2})^{1/2}}\right)G,$
so that
$G=\frac{\Lambda}{(1+X^{2})^{1/4}\sqrt{1-\sqrt{1+X^{2}}}}.$
Thus this part of $A_{n}$ is
$A_{n}\sim\frac{\Lambda}{(1+X^{2})^{1/4}\sqrt{1-\sqrt{1+X^{2}}}}\frac{(-1)^{n}3^{n}\Gamma(n+\gamma)}{(1+X^{2})^{3(n+\gamma)/2}}.$
As $X\rightarrow\pm\mathrm{i}$,
$A_{n}\sim\frac{\Lambda}{(\pm 2\mathrm{i})^{1/4}(X\mp\mathrm{i})^{1/4}}\frac{(-1)^{n}3^{n}\Gamma(n+\gamma)}{(\pm 2\mathrm{i})^{3(n+\gamma)/2}(X\mp\mathrm{i})^{3(n+\gamma)/2}}.$
To be consistent with $A_{0}$ requires $\gamma=0$ giving, finally
@@
$A_{n}\sim\frac{\bar{\lambda}_{n-1}}{(1+X^{2})^{1/4}\sqrt{1+\sqrt{1+X^{2}}}}\left(C-\tanh^{-1}\sqrt{1+X^{2}}\right).$
Again, the presence of a logarithm here means we need to modify slightly the large $n$ ansatz to allow for $\log n$ terms, which is the same as allowing the constant $C$ to depend on $\log n$. Doing so gives
$A_{n}=\frac{\bar{\lambda}_{n-1}}{(1+X^{2})^{1/4}\sqrt{1+\sqrt{1+X^{2}}}}\left(C_{0}\log n+C_{1}-\tanh^{-1}\sqrt{1+X^{2}}\right).$
Together
@@ 3.4 Boundary layer in the late terms near $X=0$
Here we come to the key step. The large $n$ asymptotic series for $A_{n}$ in the outer region is nonuniform, and rearranges when $X$ is small—there is another inner region near the origin. We emphasize again that this is a boundary layer not in the small-$\epsilon$ expansion of $A$, but in the large $n$-expansion of $A_{n}$. Again, this nonuniformity is evident because the asymptotic formula (30) is singular at $X=0$, while $A_{n}$ should be regular there. This time the appropriate scaling for the inner region is $X=\xi/n^{1/2}$ (so that $XA_{n}^{\prime}$ balances with $A_{n-1}^{\prime\prime}$), giving
@@
Both the particular integral and the complementary function exhibit Stokes’ phenomenon for large $\xi$, so that there will be a switch in the behaviour of the late terms depending on the argument of $\xi$, corresponding to the higher-order Stokes’ phenomenon. There is a higher-order Stokes line due to the particular integral when $\xi$ crosses the real axis, across which the coefficient of $\mathrm{e}^{-\xi^{2}}/\xi$ in the far field (i.e. in the outer limit of the inner expansion) changes. The branch cut associated with $\sqrt{\xi^{2}}$ in (37) must be chosen to line up with this higher-order Stokes line. In addition, there is a higher-order Stokes line on the imaginary axis, across which the constant in the far field changes. The branch cut associated with $\pm\mathrm{i}\pi/2$ must be chosen to align with this higher-order Stokes line. As $\xi\rightarrow\infty$ in the first quadrant,
@@
Figure 2: The ratio of the numerical value found by iterating (23)-(24) to the asymptotic prediction (41). Blue is the base series, while orange and green correspond to enhanced convergence using Richardson extrapolation on two and three terms respectively.
@@ 4 Example 3: Simplified Rossby waves
Our third example is a simplified version of the model problem for an equitorially-trapped Rossby wave considered in [10]. Consider
@@
with $\psi(0)=1$. Essentially the problem considered in [10] has the $\epsilon^{2}$ in the third term replaced with $\epsilon$. The switch to $\epsilon^{2}$ makes the inner region below more complicated, but significantly simplifies all the other regions of the analysis.
@@ 4.2 Outer region
As usual, there are two types of divergence: a factorial/power from repeated differentiation of the singularity $\log(1+X)$ in $\psi_{1}$, and a factorial/constant from $\lambda_{n}$. For the first, we use the usual ansatz
$\psi_{n}\sim\frac{G\Gamma(n+\gamma)}{\chi^{n+\gamma}}.$ (48)
At leading order in $n$ this gives
$\chi^{\prime}=-2X$
so that
$\chi=1-X^{2}$
since $\chi=0$ at X=1X=-1. At next order,
$G+XG^{\prime}$ $=$ $0$
so that
$G=\frac{\Lambda}{X}.$
Thus this part of $\psi_{n}$ satisfies
$\psi_{n}\sim\frac{\Lambda\Gamma(n+\gamma)}{X(1-X^{2})^{n+\gamma}}.$
As $X\rightarrow-1$,
$\psi_{n}\sim-\frac{\Lambda\Gamma(n+\gamma)}{2^{n+\gamma}(X+1)^{n+\gamma}}.$
Comparing powers of $X+1$ with the early terms gives $\gamma=-1$, so that
@@ 4.4 Boundary layer in the late terms near $X=0$
Both the particular integral and the complementary function exhibit Stokes’ phenomenon for large $\xi$, corresponding to the higher-order Stokes phenomenon in $\psi$. There is a higher-order Stokes line due to the particular integral when $\xi$ crosses the imaginary axis, across which the coefficient of $\mathrm{e}^{\xi^{2}}/\xi$ in the far field changes. In addition, there is a higher-order Stokes line on the real axis, across which the constant in the far field changes. For real $\xi$, as $\xi\rightarrow$-$\infty$
@@ 4.5 Matching with the outer
Figure 3: The ratio of the numerical value fourn by iterating (44)-(45) to the asymptotic prediction (55). Blue is the base series, while orange, green and red correspond to enhanced convergence using Richardson extrapolation on two, three and four terms respectively. The convergence is slower than expected because of the presence of logarithmic terms in the higher-order corrections, unaccounted for in the extrapolation.

### v2

@@ Abstract
We consider a class of linear eigenvalue problems depending on a small parameter $\epsilon$ in which the series expansion for the eigenvalue in powers of $\epsilon$ is divergent. We develop a new technique to determine the precise nature of this divergence. We illustrate the technique through its application to four examples: the anharmonic oscillator, a simplified model of equatorially-trapped Rossby waves, and two simplified models based on quasinormal modes of Reissner-Nordström-de Sitter black holes.
@@ 1 Introduction
Sometimes this series diverges, and it is of interest to determine the nature of this divergence. From a purely numerical point of view understanding the large-order behaviour tells you how many terms to take in an optimal approximation, how large the smallest achievable error is, and whether the series can be meaningfully resummed. However, as has long been known, large-order growth is controlled by non-perturbative effects that are not visible in ordinary perturbation theory. For eigenvalue problems, this can mean an exponentially small imaginary part of the eigenvalue, indicating an instability not visible in the regular perturbation expansion, or, in quantum mechanics, tunnelling between classically distinct regions, or exponentially-small level splittings. Here we provide a straightforward approach to determining the precise asymptotic behaviour of $\lambda_{n}$ for large $n$.
@@
The main technique of Bender and Wu is to analytically continue in $\epsilon$ (typically until $\epsilon$ is negative), and then solve the resulting problem by combination of Liouville-Green (WKB) and matched asymptotic approximations. This is a delicate procedure, since the goal is to identify an exponentially small component of $\lambda$ beyond-all-orders of the divergent asymptotic series. Cauchy’s integral formula is then used to determine the coefficients in the power series expansion of $\lambda$ on the positive real $\epsilon$ axis in terms of the values of $\lambda$ on either side of the negative real $\epsilon$ axis. This technique is ingenious, and has proved successful, but the details can be very complicated. The present work aims to present an alternative method, which we hope will be useful.
Intriguing also is the work of Dunne & Ünsal [8, 9], who use a uniform WKB approach for quantum mechanical systems with degenerate minima, focusing in particular on double-well and sine-Gordon potentials. They write the wave function in terms of a scaled parabolic cylinder function $\Psi(x)=D_{\nu}(u(x)/\sqrt{\epsilon})/\sqrt{u^{\prime}(x)}$, where the index $\nu$ is a parameter to be determined along with $u(x)$. Although $u(x)$ and $\lambda$ can be determined locally in terms of $\nu$ by a regular perturbation expansion, $\nu$ itself is determined by imposing a global boundary condition (a symmetry condition at the midpoint between wells). This shows $\nu$ is exponentially close to an integer, and provides an exponential correction to $\lambda$. In [8] this exponential correction is used to determine the late-order terms in the power series, in a similar way to Bender & Wu.
@@
In the applied mathematics literature an early example of such a problem occurs in the work of Boyd and Natarov [4], who consider a model problem for an equatorially-trapped Rossby wave in a shear flow in the ocean or atmosphere. There the main interest is in the imaginary part of the eigenvalue (corresponding to the growth rate of instability)—the divergent perturbation series is purely real, but there is an exponentially small imaginary part beyond all orders. In [12] this problem is attacked in almost the reverse direction to Bender and Wu—the divergent series is first found, and used to determine the exponentially small imaginary component of the eigenvalue via optimal truncation and Stokes phenomenon, rather than the other way round. Simplifying and extending the procedure from [12] forms the basis of the present work.
@@
We present our procedure through its application to four examples. Each follows the same general framework, but the final part of the analysis differs slightly in each case. We hope this will allow the interested reader to adapt the method to their own particular problem.
The general framework is:
(i) Inner region. Each problem starts with a regular perturbation expansion. Typically the coefficients in the expansion of the eigenvalue are determined by imposing a regularity condition at the origin $x=0$, and each term in the expansion is a polynomial in $x$. This region determines the eigenvalue expansion completely, but it is difficult to extract the large-order behaviour from the resulting recurrence relations.
(ii) Outer region. The regular expansion in (i) is not uniform, and rearranges for large $x$, leading to a new expansion once $x=X/\epsilon$ has been rescaled. The outer problem is a singular perturbation problem, so that this expansion diverges in the usual form of factorial/power, driven by singularities away from $X=0$. The large-order behaviour of this divergence is easy to determine by now standard methods. In addition the late terms have an independent component driven by the divergent eigenvalue expansion.
(iii) Boundary layer in the late terms near $X=0$. This is the new ingredient to the method. The key observation is that the large-order approximation of the outer expansion (ii) is also non-uniform, so that there is another inner region near $X=0$, now not in the small-$\epsilon$ expansion but in the large-order expansion. The resolution of this inner region links the two parts of the expansion in (ii), and determines the large-order behaviour of the eigenvalue.
@@ 2 Example 1: Simplified black holes
with $g(0)=1$ and $g(x)=o(\mathrm{e}^{-x})$ as $x\rightarrow$-$\infty$, where $0<\epsilon\ll 1$. This is a much-simplified version of the problem in [7] concerning quasinormal modes of Reissner-Nordström-de Sitter black holes, keeping only those ingredients necessary to illustrate the methodology; very roughly speaking $g$ is the charged scalar field perturbation, $x$ is the radial distance measured from the cosmological horizon, $\epsilon$ is the charge, and the eigenvalue $\omega$ is the frequency of the mode.
@@
The boundary conditions at $x=0$ and $x=$-$\infty$ each remove one degree of freedom, so that there is a nonzero solution only if $\omega$ takes particular values. The goal is to find the asymptotic expansion of the eigenvalues,
@@
as $\epsilon\rightarrow 0$, and in particular the form of the divergence of $\omega_{n}$ as $n\rightarrow\infty$. To simplify the presentation we will focus on the leading eigenvalue, here and in each of our examples, but of course the methodology works for any eigenvalue.
@@ 2.3 Inner region near $X=1$
$c_{n}=-(-1)^{n}\frac{(3/4)_{n-1}(5/4)_{n-1}}{16(2)_{n-1}},$
where $(x)_{n}=\Gamma(x+n)/\Gamma(x)$ is Pochammer’s symbol. Matching with (13) gives
@@ 2.4 Boundary layer in the late terms near $X=0$
is the exponential integral. Now $g_{n}$ should be regular as $\xi\rightarrow 0$. Since $\operatorname{Ei}(2\xi)\sim\log\xi$ as $\xi\rightarrow 0$, we need
@@
to remove the logarithmic singularity at $\xi=0$. Since $\operatorname{Ei}(2\xi)$ exhibits Stokes phenomenon for large $\xi$ there will be a switch in the behaviour of the late terms depending on the argument of $\xi$—this is what is known as the higher-order Stokes phenomenon, a Stokes phenomenon not in the asymptotic expansion of $g$ as a function of $\epsilon$, but in the late-term approximation of $g_{n}$ [6, 11]. There is a higher-order Stokes line when $\xi$ crosses the positive real axis, across which the constant contribution to the large-$\xi$ approximation of $H$ (i.e. in the outer limit of the inner expansion) changes. Note that there is no Stokes phenomenon associated with the particular solution $\log\xi$, so that the coefficient of $\mathrm{e}^{2\xi}/\xi$ is fixed. This will not be the case in our other examples.
@@
Figure 1: Divergence of the coefficients in the asymptotic expansion of $\omega$. (a) coefficients determined numerically from (4)-(5). The linear growth is consistent with factorial divergence. (b) The ratio of the numerical value to the asymptotic prediction (17). Blue is the base series, while orange, green and red correspond to enhanced convergence using Richardson extrapolation on two, three and four terms respectively. The black line is the asymptote, included to aid the eye. The convergence is slower than expected because of the presence of log terms in the higher-order corrections, unaccounted for in the extrapolation.
@@ 3 Example 2: Anharmonic oscillator
Having introduced the procedure with a simple model problem, we now consider the classical problem of the anharmonic oscillator [2, 3]. Much of the analysis follows the same framework, though the details of the boundary layer in the late-term approximation analogous to §2.4 are a little different.
@@ 3.2 Outer region
Substituting (27) into (26) and equating coefficients of powers of $\epsilon$ gives, following a routine calculation,
$\phi=\frac{1}{6}+\frac{X^{2}}{4}-\frac{(1+X^{2})^{3/2}}{6},\qquad A_{0}=\frac{\sqrt{2}}{(1+X^{2})^{1/4}\sqrt{1+\sqrt{1+X^{2}}}},$ (28)
where the normalisation $\sqrt{2}$ comes from matching with the inner region. In general, the equation for $A_{n}$ is
@@
As before, there are two sources of divergence in $A_{n}$: the usual factorial/power from repeated differentiation of the singularity in $A_{0}$, and a factorial/constant divergence driven by $\bar{\lambda}_{n}$. For the first, we use the usual factorial/power ansatz following the procedure in [5] to find
@@
$A_{n}\sim\frac{\bar{\lambda}_{n-1}}{(1+X^{2})^{1/4}\sqrt{1+\sqrt{1+X^{2}}}}\left(C_{0}\log n+C_{1}-\tanh^{-1}\sqrt{1+X^{2}}\right),$
where, as in §2.2, the presence of a logarithm in $X$ necessitates the inclusion of a logarithm in $n$. Together
@@ 3.4 Boundary layer in the late terms near $X=0$
Here we come to the key step. The large $n$ asymptotic series for $A_{n}$ in the outer region is nonuniform, and rearranges when $X$ is small—there is another inner region near the origin. We emphasise again that this is a boundary layer not in the small-$\epsilon$ expansion of $A$, but in the large $n$-expansion of $A_{n}$. Again, this nonuniformity is evident because the asymptotic formula (30) is singular at $X=0$, while $A_{n}$ should be regular there. This time the appropriate scaling for the inner region is $X=\xi/n^{1/2}$ (so that $XA_{n}^{\prime}$ balances with $A_{n-1}^{\prime\prime}$), giving
@@
Both the particular integral and the complementary function exhibit Stokes phenomenon for large $\xi$, so that there will be a switch in the behaviour of the late terms depending on the argument of $\xi$, corresponding to the higher-order Stokes phenomenon. There is a higher-order Stokes line due to the particular integral when $\xi$ crosses the real axis, across which the coefficient of $\mathrm{e}^{-\xi^{2}}/\xi$ in the far field (i.e. in the outer limit of the inner expansion) changes. The branch cut associated with $\sqrt{\xi^{2}}$ in (37) must be chosen to line up with this higher-order Stokes line. In addition, there is a higher-order Stokes line on the imaginary axis, across which the constant in the far field changes. The branch cut associated with $\pm\mathrm{i}\pi/2$ must be chosen to align with this higher-order Stokes line. As $\xi\rightarrow\infty$ in the first quadrant,
@@
Figure 2: The ratio of the numerical value found by iterating (23)-(24) to the asymptotic prediction (41). Blue is the base series, while orange and green correspond to enhanced convergence using Richardson extrapolation on two and three terms respectively. The black line is the asymptote, included to aid the eye.
@@ 4 Example 3: Simplified Rossby waves
Our third example is a simplified version of the model problem for an equatorially-trapped Rossby wave considered in [12]. Consider
@@
with $\psi(0)=1$. Essentially the problem considered in [12] has the $\epsilon^{2}$ in the third term replaced with $\epsilon$. The switch to $\epsilon^{2}$ makes the inner region below more complicated, but significantly simplifies all the other regions of the analysis. This weakens the strength of the pole in Boyd and Natarov’s Hermite-with-pole equation [4], while keeping the same structure.
@@ 4.2 Outer region
As usual, there are two types of divergence: a factorial/power from repeated differentiation of the singularity $\log(1+X)$ in $\psi_{1}$, and a factorial/constant from $\lambda_{n}$. For the first, we use the usual ansatz to give
@@ 4.4 Boundary layer in the late terms near $X=0$
Both the particular integral and the complementary function exhibit Stokes phenomenon for large $\xi$, corresponding to the higher-order Stokes phenomenon in $\psi$. There is a higher-order Stokes line due to the particular integral when $\xi$ crosses the imaginary axis, across which the coefficient of $\mathrm{e}^{\xi^{2}}/\xi$ in the far field changes. In addition, there is a higher-order Stokes line on the real axis, across which the constant in the far field changes. For real $\xi$, as $\xi\rightarrow$-$\infty$
@@ 4.5 Matching with the outer
Figure 3: The ratio of the numerical value found by iterating (44)-(45) to the asymptotic prediction (54). Blue is the base series, while orange, green and red correspond to enhanced convergence using Richardson extrapolation on two, three and four terms respectively. The black line is the asymptote, included to aid the eye. The convergence is slower than expected because of the presence of logarithmic terms in the higher-order corrections, unaccounted for in the extrapolation.